A mother is twice as old as her daughter now, but 10 years ago, she was 3 times as
old as her daughter. Find their ages at present.

Found 3 solutions by mananth, ikleyn, josgarithmetic:
Answer by mananth(16949)

Let the daughter's age be x
mother's age will be 2x
Ten years ago
daughter's age x-10
Mother's age 3(x-10)
x-10 =3(x=10)
x=10=3x-30
3x-x=10+30
2x=40
x=40/2
x=20 Daughter's age now
Mother's age is 2x = 2*20 = 40 years

Answer by ikleyn(53751)

.
A mother is twice as old as her daughter now, but 10 years ago, she was 3 times as
old as her daughter. Find their ages at present.
~~~~~~~~~~~~~~~~~~~~~~~~~~~~


        In the post by @mananth, the setup equation is written incorrectly.
        His manipulations that follow this incorrect equation, also have many errors,
        although the final answer is correct.

        His solution can not be considered as an etalon, so I came to bring a correct solution.


Let the daughter's age be x
mother's age will be 2x
Ten years ago
daughter's age x-10
Mother's age 2x-10

3(x-10) = 2x-10         <<<---=== This is a  CORRECT  setup equation

Simplify and find 'x'

3x-30 = 2x-10
3x-2x = 30-10
x = 20 Daughter's age now
Mother's age is 2x = 2*20 = 40 years

Solved correctly.



Answer by josgarithmetic(39792)